What is the energy(kJ/mol) associated with the de exitation of electron from n=6 to n=2 in He+ ion?
(a)1.36*106
(b)1.36*103
(c)1.16*103
(d)1.78*103

Grade:12

1 Answers

Vikas TU
7 years ago
For H-atom,
E2 = -13.6*4/4 = -13.6
E6 = -13.6*4/36 = -13.6/9
Excitation energy,
E2 – E6 = E
subtract them and get the energy value.